# Process-based Inference for Spatial Energetics Using Bayesian Predictive Stacking

### v2

@@ 6. Simulation
We illustrate the implementation and inferential effectiveness of our proposed methods through numerical experiments. In Section 6.1, we explore the in-fill paradigm prediction on a continuous trajectory, while Section 6.2 illustrates the performance of our proposed methods.
@@ 6.1. In-fill prediction
We consider a single trajectory $\gamma:\mathcal{T}\to\mathbb{R}^{2}$, where $\gamma(t)$ signifies the correspondence between the continuous closed interval $\mathcal{T}$ and a curve in $\mathbb{R}^{2}$. Because data are observed at discrete points on this trajectory, the accuracy of in-fill prediction is expected to improve as the number of discrete observed points grows. We consider interpolation of the outcome over a space-time trajectory composed of line segments. As an example of in-fill prediction, the conceptual diagram in Figure 1 depicts $T=20$ observations in the interval $\mathcal{T}=[1,20]$. The left panel displays the locations at each time point with the line segments comprising the trajectory shown by a green line. The right panel illustrates the space-time coordinates where we seek to predict the outcome.
We generate the data along a trajectory, which comprises our fixed domain, in accordance with the model in (4)–(7). We generate $300$ points on the trajectory over $t\in\mathcal{T}=[1,300]$ using $\gamma(t)=\gamma(t-1)+\mathcal{WN}(0,1)$, where $\mathcal{WN}(0,1)$ denotes white noise with zero mean and unit variance. We randomly include $n$ space-time coordinates over the trajectory and generate $n$ values of $z(\gamma(t),t)$ from the Gaussian process with 00 mean and covariance kernel as (5). The true parameters defining the spatial-temporal process $z(\gamma(t),t)$ in (5) are $\phi_{1}=1/2$ and $\phi_{2}=1/2$. We then generate $y(\gamma(t),t)$ from (4) using elements of $x(\gamma(t),t)$ generated from $N(0,4)$ and $\beta_{j}(t)$ using a zero-centered Gaussian process with covariance kernel in (7) specified by $\xi=1/2$. Additionally, we set $\sigma=1$, $\delta_{\beta}=1$, and $\delta_{z}=1$.
@@
6.2. Estimation performances of proposed models
We conducted simulation experiments to assess both discrete- and continuous-time trajectory models, comparing their performance in terms of estimation errors and model fitting. First, we sampled data along with the discrete time trajectory model in (2)–(3). We generated $n=T=50$ and $70$ points for two experiments on a trajectory using the same random walk model for $\gamma(t)$ as in Section 6.1. We take $p=2$ in (2) and generate initial values for the elements of ${\beta}_{0}$ and ${z}_{0}$ from $N({0},4)$. With these fixed initial values, we sequentially generate ${\beta}_{t}$ and ${z}_{t}$ using the autoregressive specification (see Section 2.2) with $\delta_{\beta}=1$, $\delta_{z}=1$, $\sigma=1$ and with $K_{\phi}$ taken as the Matérn kernel, introduced in Section 5.4, with $\phi=1/7$ and $\nu=1$. Each element of ${x}_{t}(\gamma(t))$ is generated from $N(0,4)$ and fixed thereafter. Then, $y_{t}(\gamma(t))$ is generated using (2).
We generated $30$ different datasets through the above procedure and analyzed each of them using the models in (3) and (8). For predictive stacking, we set candidate parameters in (3) as $\phi\in\{1,1/10\}$, $\nu\in\{3,1/3\}$ and $\{5,1/5\}$ for both $\delta_{\beta}$ and $\delta_{z}$. For (8), we consider $\phi_{i}\in\{3,1/10\}$ for $i=1,2$, $\xi\in\{3,1/10\}$, and $\{3,1/3\}$ for $\delta_{\beta}$ and $\delta_{z}$. We employed $20$-fold expanding window cross-validation and selected the stacking weights described in Sections 4.1 and 4.2. Additionally, we estimated a non-spatial dynamic linear model (NSDLM),
@@ 6.2. Estimation performances of proposed models
We use MSE to compare the posterior predictive means for $y_{T+1}$ with the underlying signals in (2). Similarly, we use relative mean squared errors (rMSE) defined as $\sum_{t=1}^{T}(\widehat{\omega}_{t}-\omega_{t})^{2}/(\sum_{t=1}^{T}\omega_{t}^{2})$, where $\widehat{\omega}_{t}$ denotes the posterior mean and $\omega_{t}$ is the true value, to compare our inferential effectiveness, We apply rMSE to each element of $\beta_{t}$ and $z_{t}$, while we use $(\widehat{\sigma}-\sigma)^{2}/\sigma^{2}$ for $\sigma$. To evaluate model fit, we report MLPD, the deviance information criterion (DIC, Spiegelhalter et al., 2002) and the widely applicable information criterion (WAIC, Watanabe, 2010).
@@
Figure 3 presents the posterior means (solid curve) for $y_{t}(\gamma(t))$, $z_{t}(\gamma(t))$ and the two regression coefficients comprising each $\beta_{t}$ as a function of time for one representative dataset with $n=50$ in the simulation experiment. Also shown are the true values generating the data in the form of a dashed line. We see that the posterior bands from the discrete model are more effective in containing the dashed line (truth) with the continuous model and the NSDLM still performing fairly well, although both prominently miss the truth for the second coefficient in $\beta$.
Next, we generated $30$ datasets from the continuous-time trajectory model described in (4)–(7). We generate $n=T=50$ and $70$ points on the trajectory by the same random walk as in Section 6.1. The true parameters for spatial-temporal process $z(\gamma(t),t)$ in (5) are $\phi_{1}=1/2,\phi_{2}=1/2$ and $\delta_{z}=1$. Then, we produce $y$ from (4) using $\sigma=1$, each $\beta_{j}$ generated from (7) with $\xi=1/2$ and $\delta_{\beta}=1$ and the elements of $x(\gamma(t),t)$ generated from $N(0,4)$.
We analyzed each of the $30$ datasets using the continuous time and discrete time trajectory models. For predictive stacking, a set of candidate parameters of the discrete model in (3) was set to $\phi\in\{2,1/2\}$, $\nu\in\{2,1/2\}$, $\delta_{\beta}\in\{1/2,1/10\}$ and $\delta_{z}\in\{1/2,1/10\}$. For the continuous model in (8), we set $\phi_{i}\in\{1,1/4\}$ for $i=1,2$, $\xi\in\{1,1/4\}$, and $\delta_{\beta}$ and $\delta_{z}$ in $\{3,1/3\}$. We employed $20$-fold expanding window cross-validation to determine the stacking weights as in the earlier example.
@@
Figure 4. Posterior distributions of parameters obtained by discrete time, continuous time and NSDLM models for one representative dataset generated from a continuous time trajectory model. The top panels display histograms of the posterior distribution of $\sigma$. The others show posterior summaries for $y$, $z$ and $\beta$, with solid blue lines representing posterior means, shaded blue areas representing 95% credible intervals, and dashed lines representing true values.
Figure 4 is the analogue of Figure 3 for a representative dataset from the continuous time experiment showing posterior mean and 95% posterior interval for $y(\gamma(t),t)$, $z(\gamma(t),t)$ and two slopes as a function of time. While all three methods seem comparable in their ability to capture the truth (dashed line), the precision for the continuous model is higher than the other two models with NSDLM clearly having the widest bands for the regression slopes.
@@ 7. Application
We apply the continuous space-time model (4) to the actigraph dataset, sourced from the Physical Activity through Sustainable Transport Approaches in Los Angeles (PASTA-LA), and described in detail in Alaimo Di Loro et al. (2023). Actigraph data are collected through wearable devices including sensors and a smartphone application, providing high-resolution and repeatable measurements for monitoring human activity. There is a growing body of research on the statistical relationships with physical activity measures, such as energy expenditure measures (EE) (Crouter et al., 2006; Freedson et al., 2012; Taraldsen et al., 2012) and the Metabolic Equivalent of Task (MET) (Ishikawa-Takata et al., 2008; Lyden et al., 2014; Staudenmayer et al., 2015; Bai et al., 2018). Here, we consider the instantaneous body vector Magnitude of Acceleration (MAG) as the primary endpoint of our analysis (Doherty et al., 2017). Further discussion about the conversion of MAG into energy expenditure measures is reported in (Alaimo Di Loro et al., 2023).
A major aim of this study is to estimate an individual’s MAG along a traversed path after accounting for the impact of explanatory variables (e.g., environmental features, risk factors) on metabolism. These estimates from the continuous time model, which represent statistical learning of an individual’s physical activity profile on any given day, are then used to predict the subject’s metabolic activity on any arbitrary trajectory and comprise a personalized recommendation system for the subject. The models we have developed here can yield more effective pathways and environments to enhance physical activity levels of subjects and lead to overall improvements in metabolic levels.
@@
Focusing on one individual, which is often the goal in personalized health science research seeking data driven recommendation systems for metabolic activity, we consider recordings on the MAG observed at $650$ unevenly spaced time points. The left panel of Figure 5 displays the values of the recorded MAG and the right panel plots the MAGs over the path traversed by this individual. We fit the continuous time model in (4) using “Slope”, representing the gradient at the spatial location, and “NDVI”, representing normalized vegetation index, which is a measure of greenness at the location, as the two explanatory variables in $x(\gamma(t),t)$, where $\gamma(t)$ is recorded as the subject’s coordinates along the trajectory using information from GPS. For predictive stacking, the candidates for the parameters $\phi_{1},\phi_{2},\xi$ were set $\{100,1\}$ and those for $\delta_{\beta},\delta_{z}$ were set $\{20,1/5\}$. Of the $650$ observations, $130$ data were randomly selected and used as test data, and the remaining $520$ data were utilized as training data. We computed the posterior distributions for both the continuous time trajectory model and Bayesian linear regression on the training data and performed predictions. For comparison, we applied a Bayesian linear regression model to this data using the “brm” function from the (“brms” package Bürkner, 2023) in R and adopted default priors, that is, $t_{3}(0,2.5)$ prior for $\sigma$ and flat priors for coefficients of the explanatory variables.
@@
Table 4 indicates the superiority of the proposed methods over Bayesian linear regression. Notably, there is a significant difference in MSPE with the proposed model exhibiting considerably lower errors compared with Bayesian linear regression. The MLPD, indicating the goodness of the posterior distribution, also demonstrates that the proposed model outperforms Bayesian linear regression. This superior performance is likely attributable to accounting for the spatial-temporal structure in the data. The incorporation of spatial and temporal structure into a model allows for a more detailed depiction of the data, resulting in more precise predictions. By contrast, Bayesian linear regression does not account for the information, confirming its limitations for real data where the spatial-temporal information is significant.
@@
Figure 7. Left: Posterior predictive bands of $y$ obtained by continuous time trajectory model. Right: Posterior predictive bands of $z$ obtained by continuous time trajectory model.
@@
Given the complications associated with streaming measurements at high resolutions from wirelessly operating wearable devices, it is customary to encounter swaths of time intervals that have not recorded measurements either due to technical malfunction or user behavior. Our model based inferential framework uses the posterior predictive distributions to impute such missing values. The left panel in Figure 8 presents the reconstructed MAG for the subject under consideration at some epochs using the continuous time model. The right panel presents posterior predictive credible intervals that reveal the model’s ability to effectively capture 130 held-out values for predictive validation.
@@
Finally, we compare the continuous time, discrete time and NSDLM using DIC and WAIC. For this, we extract $150$ distinct time points from the above data, ensuring that they are equally spaced and compatible with the discrete-time trajectory model in (2) and the NSDLM introduced in Section 6.2. For stacking of (2), the candidates for the parameters were $\phi\in\{1,1/10\}$, $\nu\in\{1,1/3\}$ and $\delta_{\beta},\delta_{z}\in\{5,1/5\}$. Table 5 reveals that the continuous time trajectory model is preferred (lower values) to the others in either of these metrics, while the discrete time model considerably outperforms NSDLM. In designing a physical activity recommending system that trains one model, our analysis suggests using the continuous time model is preferable although the discrete time model should also be competitive.
@@ Acknowledgments
$y_{t}(s)$ $=z_{t}(s)+\eta_{1},\quad\eta_{1}\overset{\mathrm{i.i.d.}}{\sim}N(0,\sigma^{2}),$
$z_{t}(s)$ $\overset{\mathrm{ind}}{\sim}GP\left(0,\sigma^{2}\Delta_{zt}^{2}K_{\phi}(\cdot,\cdot)\right),$
where $\Delta_{zt}^{2}=\sum_{j=1}^{t}\delta_{z}^{2j}(\alpha\sigma)^{2(j-1)}$. Applying Theorem 2.1 in Tang et al. (2021), we obtain that $\mathbb{P}_{t}^{\prime}$ and $\mathbb{P}_{t}^{*}$ are equivalent. The equivalence of the joint distributions follows. ∎

### v3

@@ 6. Simulation
We illustrate implementation and inferential effectiveness of our proposed methods with numerical experiments. Section 6.1 explores in-fill prediction on a continuous trajectory, while Section 6.2 illustrates the performance of our proposed methods.
@@ 6.1. In-fill prediction
We consider a single trajectory defined by $\gamma:\mathcal{T}\to\mathbb{R}^{2}$, where $\gamma(t)$ over time interval $\mathcal{T}$. Because data are observed at discrete points on this trajectory, the accuracy of in-fill prediction is expected to improve as the number of discrete observed points increases. We consider interpolation of the outcome over a space-time trajectory composed of line segments. As an example of in-fill prediction, the conceptual diagram in Figure 1 depicts $T=20$ observations in the interval $\mathcal{T}=[1,20]$. The left panel displays the locations at each time point with the line segments comprising the trajectory shown by a green line. The right panel illustrates the space-time coordinates where we seek to predict the outcome.
We generate data along a trajectory comprising a fixed domain in accordance with (4)–(7). We generate $300$ points on the trajectory over $t\in\mathcal{T}=[1,300]$ using $\gamma(t)=\gamma(t-1)+\mathcal{WN}(0,1)$, where $\mathcal{WN}(0,1)$ denotes white noise with zero mean and unit variance. We randomly include $n$ space-time coordinates over the trajectory and generate $n$ values of $z(\gamma(t),t)$ from the Gaussian process with 00 mean and covariance kernel as (5). The parameters defining the spatial-temporal process $z(\gamma(t),t)$ in (5) are set to $\phi_{1}=1/2$ and $\phi_{2}=1/2$. We then generate $y(\gamma(t),t)$ from (4) using elements of $x(\gamma(t),t)$ generated from $N(0,4)$ and $\beta_{j}(t)$ using a zero-centered Gaussian process with covariance kernel in (7) specified by $\xi=1/2$. Additionally, we set $\sigma=1$, $\delta_{\beta}=1$, and $\delta_{z}=1$.
@@
6.2. Estimation of proposed models
We conduct simulation experiments to assess discrete- and continuous-time trajectory models, comparing their performance in terms of estimation errors and model fitting. First, we sample data using the discrete time trajectory in (2)–(3). We generate $n=T=50$ and $70$ points for two experiments on a trajectory using the same random walk model for $\gamma(t)$ as in Section 6.1. We let $p=2$ in (2) and generate values for the elements of ${\beta}_{0}$ and ${z}_{0}$ from $N({0},4)$. With these fixed values, we sequentially generate ${\beta}_{t}$ and ${z}_{t}$ using the autoregressive model (Section 2.2) with $\delta_{\beta}=1$, $\delta_{z}=1$, $\sigma=1$ and $K_{\phi}$ taken as the Matérn kernel introduced in Section 5.4 with $\phi=1/7$ and $\nu=1$. Each element of ${x}_{t}(\gamma(t))$ is generated from $N(0,4)$ and fixed thereafter. Then, $y_{t}(\gamma(t))$ is generated using (2).
We generate $30$ different datasets and analyze each of them using the models in (3) and (8). For predictive stacking, we set candidate parameters in (3) as $\phi\in\{1,1/10\}$, $\nu\in\{3,1/3\}$ and $\{5,1/5\}$ for both $\delta_{\beta}$ and $\delta_{z}$. For (8), we consider $\phi_{i}\in\{3,1/10\}$ for $i=1,2$, $\xi\in\{3,1/10\}$, and $\{3,1/3\}$ for $\delta_{\beta}$ and $\delta_{z}$. We employ $20$-fold expanding window cross-validation and select the stacking weights described in Sections 4.1 and 4.2. We also estimate a non-spatial dynamic linear model (NSDLM),
@@ 6.2. Estimation of proposed models
We use MSE to compare posterior predictive means for $y_{T+1}$ with underlying signals in (2). We use relative mean squared errors (rMSE) defined as $\sum_{t=1}^{T}(\widehat{\omega}_{t}-\omega_{t})^{2}/(\sum_{t=1}^{T}\omega_{t}^{2})$, where $\widehat{\omega}_{t}$ is the posterior mean and $\omega_{t}$ is the true value, to compare our inferential effectiveness, We apply rMSE to each element of $\beta_{t}$ and $z_{t}$, while we use $(\widehat{\sigma}-\sigma)^{2}/\sigma^{2}$ for $\sigma$. To evaluate model fit, we report MLPD, the deviance information criterion (DIC, Spiegelhalter et al., 2002) and the widely applicable information criterion (WAIC, Watanabe, 2010).
[…]
@@
Figure 3 presents the posterior means (solid curve) for $y_{t}(\gamma(t))$, $z_{t}(\gamma(t))$ and two regression coefficients in each $\beta_{t}$ as a function of time for one representative dataset with $n=50$ in the simulation experiment. Also shown are the true values generating the data in the form of a dashed line. The posterior bands from the discrete model are effective in containing the dashed line (truth) with the continuous model and NSDLM still performing fairly well but missing the truth for the second coefficient in $\beta$.
Next, we generate $30$ datasets from the continuous-time trajectory model described in (4)–(7). We generate $n=T=50$ and $70$ points on the trajectory by the same random walk as in Section 6.1. The true parameters for spatial-temporal process $z(\gamma(t),t)$ in (5) are $\phi_{1}=1/2,\phi_{2}=1/2$ and $\delta_{z}=1$. Then, we produce $y$ from (4) using $\sigma=1$, each $\beta_{j}$ generated from (7) with $\xi=1/2$, $\delta_{\beta}=1$ and elements of $x(\gamma(t),t)$ generated from $N(0,4)$.
We analyze each of the $30$ datasets using continuous time and discrete time trajectory models. For predictive stacking, we use $\phi\in\{2,1/2\}$, $\nu\in\{2,1/2\}$, $\delta_{\beta}\in\{1/2,1/10\}$ and $\delta_{z}\in\{1/2,1/10\}$ for the discrete model in (3) . For the continuous model in (8), we set $\phi_{i}\in\{1,1/4\}$ for $i=1,2$, $\xi\in\{1,1/4\}$, and $\delta_{\beta}$ and $\delta_{z}$ in $\{3,1/3\}$. We employ $20$-fold expanding window cross-validation to determine the stacking weights as earlier.
@@
Figure 4. Posterior distributions of parameters from discrete time, continuous time and NSDLM models for one representative dataset generated from a continuous time trajectory model. Solid blue lines are posterior means, shaded blue areas represent 95% credible intervals, and dashed lines represent true values.
Figure 4 is analogous to Figure 3 for a representative dataset from the continuous time experiment showing posterior mean and 95% intervals for $y(\gamma(t),t)$, $z(\gamma(t),t)$ and two slopes as a function of time. While all three methods seem comparable in capturing the truth (dashed line), the precision for the continuous model is higher than the other two models with NSDLM clearly having the widest bands for the regression slopes.
@@ 7. Application
We apply (4) to an actigraph dataset from the Physical Activity through Sustainable Transport Approaches in Los Angeles (PASTA-LA) described in Alaimo Di Loro et al. (2023). Actigraph data are collected through wearable devices including sensors and a smartphone application, providing high-resolution and repeatable measurements for monitoring human activity. There is a growing body of research on relationships between Magnitude of Acceleration (MAG) (Ishikawa-Takata et al., 2008; Lyden et al., 2014; Staudenmayer et al., 2015; Bai et al., 2018) and physical activity measures such as energy expenditure (EE) (Crouter et al., 2006b; Freedson et al., 2012; Taraldsen et al., 2012). We consider instantaneous body vector MAG as the primary endpoint (Doherty et al., 2017).
We seek to estimate an individual’s MAG along a traversed path after accounting for the impact of explanatory variables on metabolism. These estimates from the continuous time model, which represent statistical learning of an individual’s physical activity profile on any given day, are then used to predict the subject’s metabolic activity on any arbitrary trajectory and comprise a personalized recommendation system for the subject. The models we have developed here can yield more effective pathways and environments to enhance physical activity levels of subjects and lead to overall improvements in metabolic levels.
@@
Focusing on one individual, which is often the goal in personalized health science research seeking data driven recommendation systems for metabolic activity, we consider MAGs observed at $650$ unevenly spaced time points. The left panel of Figure 5 displays the MAG values and the right panel plots the MAGs over the path traversed by this individual. We fit the model in (4) using “Slope”, representing the gradient at the spatial location, and “NDVI”, representing normalized vegetation index, which is a measure of greenness at the location, as the two explanatory variables in $x(\gamma(t),t)$. For predictive stacking, the candidates for the parameters $\phi_{1},\phi_{2},\xi$ are set $\{100,1\}$ and those for $\delta_{\beta},\delta_{z}$ are set $\{20,1/5\}$. Of the $650$ observations, $130$ are randomly selected and used as test data, and the remaining $520$ are utilized as training data. We compute the posterior and predictive distributions for the continuous time trajectory model and Bayesian linear regression on the training data. For comparison, we apply a Bayesian linear regression model to this data using the “brm” function from the (“brms” package Bürkner, 2023) in R and adopt a (default) $t_{3}(0,2.5)$ prior for $\sigma$ and flat priors for coefficients of the explanatory variables.
@@
Table 4 indicates the superiority of the proposed methods over Bayesian linear regression. There is significant difference in MSPE with the proposed model exhibiting considerably lower errors compared with Bayesian linear regression. The MLPD, indicating the goodness of the posterior distribution, unsurprisingly, demonstrates that the proposed model outperforms Bayesian linear regression by accounting for spatial-temporal structure in the data. The benefits of driving the inference using a spatial-temporal process are clear from this analysis. In particular, linear regression and ordinary least squares estimates. while informing about the impact of explanatory variables on MAGs, is severely limited in predictive capabilities as it fails to account for spatial-temporal information.
@@
Figure 7. Predictive bands of $y$ (left) and $z$ (right) from continuous time trajectory model.
@@
Given the complications associated with streaming measurements at high resolutions from wearable devices, it is customary to encounter swaths of time intervals that have not recorded measurements either due to technical malfunction or user behavior. Our model based inferential framework uses the posterior predictive distributions to impute such missing values. The left panel in Figure 8 presents the reconstructed MAG for the subject under consideration at some epochs using the continuous time model. The right panel presents posterior predictive credible intervals that reveal the model’s ability to effectively capture 130 held-out values for predictive validation.
@@
Finally, we compare the continuous time, discrete time and NSDLM using DIC and WAIC. We extract $150$ distinct time points from the above data, ensuring that they are equally spaced and compatible with the discrete-time trajectory model in (2) and the NSDLM introduced in Section 6.2. For stacking of (2), the candidates for the parameters were $\phi\in\{1,1/10\}$, $\nu\in\{1,1/3\}$ and $\delta_{\beta},\delta_{z}\in\{5,1/5\}$. Table 5 reveals that the continuous time trajectory model is preferred (lower values) to the others in either of these metrics, while the discrete time model considerably outperforms NSDLM. In designing a physical activity recommending system that trains one model, our analysis suggests using the continuous time model is preferable although the discrete time model should also be competitive.
@@ Acknowledgments
$y_{t}(s)$ $=z_{t}(s)+\eta_{1t}(s),\quad z_{t}(s)\overset{\mathrm{ind}}{\sim}GP\left(0,\sigma^{2}\Delta_{zt}^{2}K_{\phi}(\cdot,\cdot)\right),$
where $\eta_{1t}=(\eta_{1t}(s_{1}),\ldots,\eta_{1t}(s_{n}))^{\mathrm{\scriptscriptstyle T}}\overset{\mathrm{i.i.d.}}{\sim}N(0,\sigma^{2}I_{n})$ over any finite set of $n$ spatial locations within a bounded region, and $\Delta_{zt}^{2}=\sum_{j=1}^{t}\delta_{z}^{2j}(\alpha\sigma)^{2(j-1)}$. Applying Theorem 2.1 in Tang et al. (2021), we obtain that $\mathbb{P}_{t}^{\prime}$ and $\mathbb{P}_{t}^{*}$ are equivalent. The equivalence of the joint distributions follows. ∎